# Weylian.bib

### v1

@@ Abstract
Second-order (maximally) conformally superintegrable Hamiltonian systems of abundant type are re-examined, revealing their underlying Weylian geometry. This allows one to naturally extend the concept of c-superintegrability from the realm of conformal geometries to that of Weylian manifolds. The resulting structures, in any dimension, are characterised by conformally invariant equations.
@@ 1 Introduction
A geometric characterisation of these systems has recently been put forward [KSV2023, KSV2024, KSV2024_bauhaus], proposing a tensorial formalism that makes higher dimensions accessible to an efficient investigation. It revealed that the Riemannian metrics underlying such superintegrable systems are often Hessian metrics [AV2024] and that examples on constant curvature spaces arise from solutions to the Witten-Dijkgraaf-Verlinde-Verlinde (WDVV) equation [KSV2024]. Stäckel transformations, also known as coupling constant metamorphosis, are a well-known transformation of second-order superintegrable systems, which naturally embeds into the tensorial framework mentioned earlier [KSV2024, KSV2024_bauhaus]. The first of these references argues that these systems should be considered as structures on conformal manifolds (called c-superintegrability therein). The purpose of the present paper is to clarify this concept by shedding light onto it from the angle of Weylian geometry. As a result, we extend the definition of abundant superintegrable systems to Weylian structures, including the case of $2$-dimensional ones (while the ideas of [KSV2024] are general, there is a focus on $n\geq 3$ regarding the abundant structural equations and c-superintegrability). As we are going to see, c.f. Definitions 4 and 5, subtleties arise if the underlying Weylian geometry is fixed in advance. In contrast, the abundant structure itself defines a preferred Weylian structure, which may differ from a given one. We characterise this difference by invariant data. In Section 2.1, we review and improve the invariant structural equations for dimensions $n\geq 3$, c.f. [KSV2024]. In Section 2.1, we then conclude the paper by determining the analogous invariant equations for the structural equations in dimension $n=2$, based on the structural equations given in [KSV2024_bauhaus], which differ significantly from those in higher dimensions.
@@ 1.1 Weylian manifolds
The purpose of the present section is to introduce Weylian manifolds, also known as Weyl manifolds, e.g. [Folland1970, Norden1976, CP1999, BFM2023]. Following [MS2020], we use the attribute Weylian in order to avoid confusion with the concept of Weyl manifolds in Cartan geometry, c.f. [CM2023, CS2009]. The following twofold definition is based on the one given in [MS2020], see also the references therein.
@@ 1.2 Abundant superintegrable systems
1.3 Abundant systems
In the present paper, we also need the concept of abundant systems. For the following definition, we introduce the $1$-form $\beta\in\Omega^{1}(M)$,
$\beta_{k}$ $=S_{kab}\Xi^{ab}\,.$
@@ 1.3 Abundant systems
Definition 2 ([KSV2024]).
@@
    00 0 $=\Pi_{\circ}\left(g^{ab}S_{aik}S_{bjl}-g^{ab}S_{ail}S_{bjk}\right)$ (5c)
@@
    We then say that $(S,t)$ is an abundant system on $(M,g)$. If we consider only the underlying smooth manifold $M$ to be fixed, we call $(g,S,t)$ an abundant system on $M$.
    The relationship between abundant systems and superintegrable systems has been elucidated by [KSV2024, KSV2024_bauhaus], and we refer the interested reader to these references for more details: Here we confine ourselves to mentioning that given an abundant manifold, one may integrate a certain system of partial differential equations whose solutions solve (3). In fact, one typically obtains $\frac{1}{2}(n-1)(n+2)$ linearly independent functions $F^{(\alpha)}$ of the form (2) that satisfy (3) for a non-degenerate multi-Hamiltonian. The interested reader will find more details on this construction in [KSV2024, KSV2024_bauhaus] and [KKM2018] as well as in the references therein. In particular, it is shown in [KSV2024] that the equations in Definition 2 are conformally invariant. By this we mean that these equations remain true if we replace $g\mapsto\Omega^{2}g$ for some nowhere vanishing $\Omega\in\mathcal{C}^{\infty}(M)$ (and its Levi-Civita connection accordingly) as well as
@@
    Two abundant systems $(g,S,t)$ and $(g^{\prime},S^{\prime},t^{\prime})$ on the same smooth manifold are said to be conformally related if $g^{\prime}=\Omega^{2}g$, $S^{\prime}=\Omega^{2}S$ and $t^{\prime}=t-3\ln|\Omega|+c$ for some $c\in\mathbb{R}$.
    2 Abundant systems on Weylian manifolds
    Having introduced abundant systems and Weylian manifolds, we are now going to relate these two structures. Recall that for an exact Weyl structure $(M,c,D)$, the $1$-form $\theta_{g}$ in (1) is exact for any $g\in c$. We thus have $\theta_{g}=d\Theta_{g}$ for some function $\Theta_{g}\in\mathcal{C}^{\infty}(M)$. Analogously, consider the corresponding Weylian structure $(M,\Phi)$. For $(g,\varphi)\in\Phi$, the $1$-form $\varphi$ is also exact, and $\varphi=2d\phi$ for some function $\phi$, c.f. Remark 1. With this terminology, we state the following lemma, which allows us to define abundant systems on Weylian manifolds. Its proof is implied by [KSV2024].
@@ 2 Abundant systems on Weylian manifolds
    Let $(M,\Phi)$ be an exact Weylian structure. Let $(g,\varphi)\in\Phi$ and assume that $(M,g)$ admits an abundant system encoded in the tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ and the function $t\in\mathcal{C}^{\infty}(M)$. Next, let $(g^{\prime},\varphi^{\prime})$ be another representative of the Weylian metric $\Phi$.
    Then $(M,g^{\prime})$ carries the abundant system encoded in the structure tensor $S^{\prime}=g^{\prime}g^{-1}S$ and the function $t^{\prime}=t-3(\varphi-\varphi^{\prime})$.
@@
    Clearly, $(g^{\prime},\varphi^{\prime})=(\Omega^{2}g,\varphi-\ln|\Omega|)$ for some function $\Omega$. We hence have $S^{\prime}=\Omega^{2}S$ and $t^{\prime}=t-3\ln|\Omega|$. It then follows from [KSV2024] that $(M,g^{\prime},S^{\prime},t^{\prime})$ is an abundant system, given that $(M,g,S,t)$ is one, and since all objects satisfy the properties of conformally related systems. The claim follows. ∎
    We observe that $\hat{S}=g^{-1}S\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ and $\hat{S}=(g^{\prime})^{-1}S^{\prime}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ coincide. We also note that in the lemma, $\frac{1}{3}t$ and $\varphi$ transform into $\frac{1}{3}t^{\prime}$ and $\varphi^{\prime}$, respectively, in an analogous manner. The equation $t-3\varphi=t^{\prime}-3\varphi^{\prime}$ hence yields a conformal invariant. We are now ready to define abundant systems on Weylian manifolds.
@@
    Compare Definition 4 to the definition of c-superintegrable systems, c.f. Definition 3.9 in [KSV2024]. We note that our focus here is on the case of abundant systems, but that the underpinning ideas are nontheless of a general nature. Indeed, we observe the fundamental difference that, in the reference, a c-superintegrable system is defined on a merely conformal structure $(M,c)$, i.e. a smooth manifold with a conformal class of metrics. The discussion in the current section has revealed the underpinning Weylian geometry. In [KSV2023], this underlying Weylian structure exists implicitly, c.f. Section 6 of [KSV2024]. For $n\geq 3$, it is indirectly inferred, by Eq. (6.6) of [KSV2023], from a conformal class of abundant superintegrable systems.
    We may reinterpret this as having subsumed the invariant difference function $\mathfrak{t}=t-3\varphi$ bridging between the representative $(g,\varphi)\in\Phi$ of the Weylian metric and the abundant system $(M,g,S,t)$. Indeed, if $\mathfrak{t}\neq 0$, we may replace $\Phi$ be a new Weylian metric $\Phi^{\prime}=[(g,\varphi+\frac{1}{3}\mathfrak{t})]$. In terms of the corresponding Weyl structure this presents itself as a conformal change of the Weylian connection. This observation now allows us to extend abundant systems to pre-fixed Weylian manifolds.
    In the light of the remark, let the Weylian manifold be fixed, e.g. in terms of a Weyl structure $(M,g,D)$. We may then define an abundant system on this Weylian manifold as follows.
@@
    Let $(M,\Phi)$ be a Weylian structure. We call $(\hat{S},\mathfrak{t})$ consisting of a (2,1)-tensor field $\hat{S}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ and a function $\mathfrak{t}\in\mathcal{C}^{\infty}(M)$ an abundant system on $(M,\Phi)$ if $(M,[(g,\varphi+\frac{1}{3}\mathfrak{t})],\hat{S})$ is an abundant Weylian manifold.
@@
      An abundant system $(\hat{S},\mathfrak{t})$ on a Weylian manifold $(M,\Phi)$ is a natural abundant Weylian structure precisely if $\mathfrak{t}$ is constant.
@@
      If $(\hat{S},\mathfrak{t})$ is an abundant system on the Weylian manifold $(M,\Phi)$, then $(M,[(g,\varphi+\mathfrak{t})],\hat{S})$ is a natural abundant Weylian manifold, $(g,\varphi)\in\Phi$.
@@ 2.1 Structural equations for $n\geq 3$
    Our aim is now to give a “representative-free” characterisation of abundant systems on Weylian manifolds. To this end, we need to establish some terminology for a given Weyl structure $(M,g,D)$. In order to do this efficiently, we introduce some terminology. We denote by $\Pi_{\mathrm{Sym}^{3}_{0}}$ the projection onto the trace-free totally symmetric component, and we introduce conformally adapted operators $\mathscr{H}:\mathcal{C}^{\infty}(M)\to\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$,
@@
    An analogous statement holds for abundant systems (with non-vanishing $\mathfrak{t}$) on flat Weylian structures, the details of which we leave to the interested reader.

### v2

@@ Abstract
Abundant second-order maximally conformally superintegrable Hamiltonian systems are re-examined, revealing their underlying natural Weyl structure and offering a clearer geometric context for the study of Stäckel transformations (also known as coupling constant metamorphosis). This also allows us to naturally extend the concept of conformal superintegrability from the realm of conformal geometries to that of Weyl structures. It enables us to interpret superintegrable systems of the above type as semi-Weyl structures, a concept related to statistical manifolds and affine hypersurface theory.
@@ 1 Introduction
In recent years, considerable advances have been made in the understanding of the geometric structures underlying maximally superintegrable systems. For instance, for two-dimensional non-degenerate systems in the Euclidean plane, [Kress&Schoebel] establishes an algebraic geometric classification, associating a unique planar line triple arrangement to every superintegrable system. For three-dimensional systems, [Capel&Kress, Capel_phdthesis] provides an explicit description of the conformal classes of these systems using algebraic varieties formed from representations of the conformal group. To facilitate an extension of the classification to higher dimensions, a geometric characterisation of non-degenerate systems has recently been put forward, cf. [KSV2023, KSV2024, KSV2024_bauhaus], proposing a tensorial formalism that makes higher dimensions accessible to an efficient investigation. It revealed that the Riemannian metrics underlying such superintegrable systems are often Hessian metrics [AV2025, CV2025] and that examples on constant curvature spaces arise from solutions to the Witten-Dijkgraaf-Verlinde-Verlinde (WDVV) equation [KSV2024, Vollmer2024].
Stäckel transformations, also known as coupling constant metamorphosis, are a well-known transformation of second-order superintegrable systems, which naturally embeds into the tensorial framework mentioned earlier [KSV2024, KSV2024_bauhaus]. The first of these references argues that these systems should be considered as structures on conformal manifolds (called c-superintegrability therein).
The purpose of the present paper is to clarify the concept of c-superintegrability by shedding light onto it from the angle of Weylian geometry, closing a gap in the literature. We explain how an abundant superintegrable system precisely gives rise to a Weyl structure. We furthermore extend the definition of abundant superintegrable systems to Weylian manifolds, including the case of $2$-dimensional systems. Note that while the ideas of [KSV2024] work in general dimension, there is a focus on $n\geq 3$ regarding the abundant structural equations and c-superintegrability. As we are going to see, c.f. Definitions 4 and 5 below, subtleties arise if the underlying Weylian geometry is fixed in advance. In contrast, the abundant structure itself defines a preferred Weylian structure, which may differ from a given one. We characterise this difference by invariant data. In Section 3.1, we review and improve the invariant structural equations for dimensions $n\geq 3$, c.f. [KSV2024]. In Section 3.1, we determine the analogous invariant equations for the structural equations in dimension $n=2$, based on the structural equations given in [KSV2024_bauhaus], which differ significantly from those in higher dimensions. We conclude the paper in Section 3.3 showing that abundant superintegrable systems can be realised as semi-Weyl structures. These structures have recently received some attention in the context of statistical manifolds and affine hypersurface geometry, e.g. [BN2022, Matsuzoe2001, Matsuzoe2010].
The paper is organised as follows: In the following section, we introduce the main objects of the paper, namely Weylian manifolds and abundant superintegrable systems in subsections 2.1 and 2.2, respectively. In Section 3, the main discussion of the paper will take place. Specifically, in Section 3.1, we discuss the case of systems of dimension $n>2$, while Section 3.2 is devoted to the case of $2$-dimensional systems, for which the underpinning structural equations are considerably different from higher dimensions. Finally, we use so-called semi-Weyl structures (a generalisation of Weyl structures) to encode abundant superintegrable systems, see Section 3.3.
2 Preliminaries
@@ 2.1 Weylian manifolds
The purpose of the present section is to introduce Weylian manifolds, also known as Weyl manifolds, e.g. [Folland1970, Norden1976, CP1999, BFM2023]. Following [MS2020], we use the attribute Weylian in order to avoid confusion with the concept of Weyl manifolds in Cartan geometry, c.f. [CM2023, CS2009]. The ensuing twofold definition follows [MS2020], see also the references therein.
@@ 2.2 Abundant superintegrable systems
We now introduce the concept of abundant superintegrable systems. For the following definition, we will denote the Ricci tensor of the Riemannian metric $g$ by $\mathrm{Ric}$,
$\mathrm{Ric}(X,Y)=\mathrm{tr}(R^{g}(\cdot,X)Y),$
where $R^{g}$ is the curvature tensor of $g$. The scalar curvature of $g$ is going to be denoted by $R$ as there is no real risk of confusion with the curvature tensor. The Schouten tensor of $g$ is going to be denoted by $\mathsf{P}$, i.e.
$\mathsf{P}=\frac{1}{n-2}\left(\mathrm{Ric}-\frac{R}{2(n-1)}g\right)$
@@
The following definition is based on the structural equations established in [KSV2023, KSV2024, KSV2024_bauhaus]. The terminology follows [CV2025].
Definition 2 ([KSV2024, KSV2024_bauhaus, CV2025]).
@@
    00 0 $=\Pi_{\circ}\left(g^{ab}S_{aik}S_{bjl}-g^{ab}S_{ail}S_{bjk}\right).$ (5c)
@@
    where we introduce the $1$-form $\beta\in\Omega^{1}(M)$, defined by
    $\beta_{k}$ $=S_{kab}\Xi^{ab}\,.$
    We then say that $(S,t)$ is an abundant structure on $(M,g)$. If we consider only the underlying smooth manifold $M$ to be fixed, we call $(g,S,t)$ an abundant structure on $M$.
    The tensor fields $\Xi,\tau\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$ in Definition 2 will in the following be considered as mostly technical, auxiliary objects. However, we mention that $\tau$ has a deeper significance for distinguishing properly and conformally superintegrable systems, cf. [KSV2024, BLMS24] and also [GKL2024].
    The relationship between abundant structures and superintegrable systems has been elucidated by [KSV2024, KSV2024_bauhaus], and we refer the interested reader to these references for more details: Here we confine ourselves to mentioning that given an abundant manifold, one may integrate a certain system of partial differential equations whose solutions solve (3). In fact, one typically obtains $\frac{1}{2}(n-1)(n+2)$ linearly independent functions $F^{(\alpha)}$ of the form (2) that satisfy (3) for a non-degenerate multi-Hamiltonian. The interested reader will find more details on this construction in [KSV2024, KSV2024_bauhaus] and [KKM2018] as well as in the references therein. In particular, it is shown in [KSV2024] that the equations in Definition 2 are conformally invariant. By this we mean that these equations remain true if we replace $g\mapsto\Omega^{2}g$ for some nowhere vanishing $\Omega\in\mathcal{C}^{\infty}(M)$ (and its Levi-Civita connection accordingly) as well as
@@
    Two abundant structures $(g,S,t)$ and $(g^{\prime},S^{\prime},t^{\prime})$ on the same smooth manifold are said to be conformally related if $g^{\prime}=\Omega^{2}g$, $S^{\prime}=\Omega^{2}S$ and $t^{\prime}=t-3\ln|\Omega|+c$ for some $c\in\mathbb{R}$.
    3 Abundant structures and Weylian manifolds
    Having introduced abundant structures and Weylian manifolds, we are now going to relate these two structures. Recall that for an exact Weyl structure $(M,c,D)$, the $1$-form $\theta_{g}$ in (1) is exact for any $g\in c$. We thus have $\theta_{g}=d\Theta_{g}$ for some function $\Theta_{g}\in\mathcal{C}^{\infty}(M)$. Analogously, consider the corresponding Weylian structure $(M,\Phi)$. For $(g,\varphi)\in\Phi$, the $1$-form $\varphi$ is also exact, and $\varphi=2d\phi$ for some function $\phi$, c.f. Remark 1. With this terminology, we state the following lemma, which allows us to define abundant structures on Weylian manifolds. Its proof is implied by [KSV2024].
@@ 3 Abundant structures and Weylian manifolds
    Let $(M,\Phi)$ be an exact Weylian structure. Let $(g,\varphi)\in\Phi$ and assume that $(M,g)$ admits an abundant structure encoded in the tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ and the function $t\in\mathcal{C}^{\infty}(M)$. Next, let $(g^{\prime},\varphi^{\prime})$ be another representative of the Weylian metric $\Phi$.
    Then $(M,g^{\prime})$ carries the abundant structure encoded in the structure tensor $S^{\prime}=g^{\prime}g^{-1}S$ and the function $t^{\prime}=t-3(\varphi-\varphi^{\prime})$.
@@
    Clearly, $(g^{\prime},\varphi^{\prime})=(\Omega^{2}g,\varphi-d\ln|\Omega|)$ for some function $\Omega$. We hence have $S^{\prime}=\Omega^{2}S$ and $t^{\prime}=t-3\ln|\Omega|$. It then follows from [KSV2024] that $(M,g^{\prime},S^{\prime},t^{\prime})$ is an abundant structure, given that $(M,g,S,t)$ is one, and since all objects satisfy the properties of conformally related systems. The claim follows. ∎
    We observe that $\hat{S}=g^{-1}S\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ and $\hat{S}=(g^{\prime})^{-1}S^{\prime}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ coincide. We also note that in the lemma, $\frac{1}{3}t$ and $\varphi$ transform into $\frac{1}{3}t^{\prime}$ and $\varphi^{\prime}$, respectively, in an analogous manner. The equation $t-3\varphi=t^{\prime}-3\varphi^{\prime}$ hence yields a conformal invariant. We are now ready to define abundant structures on Weylian manifolds.
@@
    Compare Definition 4 to the definition of c-superintegrable systems, c.f. Definition 3.9 in [KSV2024]. We note that our focus here is on the case of abundant structures, but that the underpinning ideas are nontheless of a general nature. Indeed, we observe the fundamental difference that, in the reference, a c-superintegrable system is defined on a merely conformal structure $(M,c)$, i.e. a smooth manifold with a conformal class of metrics. The discussion in the current section has revealed the underpinning Weylian geometry. In [KSV2023], this underlying Weylian structure exists implicitly, c.f. Section 6 of [KSV2024]. For $n\geq 3$, it is indirectly inferred, by Eq. (6.6) of [KSV2023], from a conformal class of abundant superintegrable systems.
    We may reinterpret this as having subsumed the invariant difference function $\mathfrak{t}=t-3\varphi$ bridging between the representative $(g,\varphi)\in\Phi$ of the Weylian metric and the abundant structure $(M,g,S,t)$. Indeed, if $\mathfrak{t}\neq 0$, we may replace $\Phi$ be a new Weylian metric $\Phi^{\prime}=[(g,\varphi+\frac{1}{3}\mathfrak{t})]$. In terms of the corresponding Weyl structure this presents itself as a conformal change of the Weylian connection. This observation now allows us to extend abundant structures to pre-fixed Weylian manifolds.
    In the light of the remark, let the Weylian manifold be fixed, e.g. in terms of a Weyl structure $(M,g,D)$. We may then define an abundant structure on this Weylian manifold as follows.
@@
    Let $(M,\Phi)$ be a Weylian structure. We call $(\hat{S},\mathfrak{t})$ consisting of a (2,1)-tensor field $\hat{S}\in\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M)\otimes TM)$ and a function $\mathfrak{t}\in\mathcal{C}^{\infty}(M)$ an abundant structure on $(M,\Phi)$ if $(M,[(g,\varphi+\frac{1}{3}\mathfrak{t})],\hat{S})$ is an abundant Weylian manifold.
@@
      An abundant structure $(\hat{S},\mathfrak{t})$ on a Weylian manifold $(M,\Phi)$ is a natural abundant Weylian structure precisely if $\mathfrak{t}$ is constant.
@@
      If $(\hat{S},\mathfrak{t})$ is an abundant structure on the Weylian manifold $(M,\Phi)$, then $(M,[(g,\varphi+\mathfrak{t})],\hat{S})$ is a natural abundant Weylian manifold, $(g,\varphi)\in\Phi$.
@@ 3.1 Structural equations for $n\geq 3$
    Our aim is now to give a “representative-free” characterisation of abundant structures on Weylian manifolds. To this end, we need to establish some terminology for a given Weyl structure $(M,g,D)$. In order to do this efficiently, we introduce some terminology. We denote by $\Pi_{\mathrm{Sym}^{3}_{0}}$ the projection onto the trace-free totally symmetric component, and we introduce conformally adapted operators $\mathscr{H}:\mathcal{C}^{\infty}(M)\to\Gamma(\mathrm{Sym}^{2}_{\circ}(T^{*}M))$,
@@
    An analogous statement holds for abundant structures (with non-vanishing $\mathfrak{t}$) on flat Weylian structures, the details of which we leave to the interested reader.
@@ 3.2 Structural equations for $n=2$
    3.3 Semi-Weyl manifolds
    Re-examining the geometric structure of abundant superintegrable systems, we have explicitly described the underlying Weyl structure. This allowed us to encode conformal classes of abundant structures in a totally symmetric and trace-free tensor field $S\in\Gamma(\mathrm{Sym}^{3}_{\circ}(T^{*}M))$ on a Weyl structure $(M,c,D)$. Consider an abundant structure $(M,g,S,t)$. Then the corresponding natural Weyl structure is given by the representative $(g_{0},0)$ where $g_{0}=e^{-\frac{2}{3}t}$. Indeed, recalling Lemma 1, a representative of the Weylian metric is given by
    $\left(\Omega^{2}g,-\frac{1}{3}\,dt-d\ln|\Omega|\right)$
    where $\Omega$ is a smooth function. Asking that $dt-d\ln|\Omega|=0$ (we recall that the Weylian manifold is exact), we obtain $\Omega^{2}=\exp(-\frac{2}{3}t)$ (modulo multiplication by an irrelevant constant). It follows that, letting $g_{0}:=e^{-\frac{2}{3}t}g$,
    $(g_{0},0)\sim\left(g,-\frac{1}{3}dt\right).$
    We hence have
    $Dg_{0}=0$
    where $D$ is the Weyl connection associated to the Weylian metric $\Phi:=[(g,-\frac{1}{3}dt)]$, compare Definition 1. This shows:
    Theorem 4.
    The natural abundant Weylian structure (cf. Definition 4) of an abundant structure $(M,g,S,t)$ is precisely the Levi-Civita connection of the metric $g_{0}=\exp(-\frac{2}{3}t)g$.
    We remark that, in the terminology of [KSV2024], the metric $g_{0}$ is the metric of the “system in standard scale” associated to the system corresponding to the abundant structure $(M,g,S,t)$ of the theorem.
    Changing our vantage point, we will now explain how this allows us to conceptualise the abundant structure as a semi-Weyl manifold. We find the following definition in [Matsuzoe2010].
    Definition 6.
    A semi-Weyl structure is a quadruple $(M,g,\bar{D},\theta)$ consisting of a Riemannian manifold $(M,g)$, a torsion-free affine connection $D$ on $M$, and a $1$-form $\theta\in\Omega^{1}(M)$, such that
    $\bar{D}g+\theta\otimes g\in\Gamma(\mathrm{Sym}^{3}(T^{*}M))$
    is totally symmetric.
    Consider an abundant structure $(M,g,S,t)$ as before and set $g_{0}=\exp(-\frac{2}{3}\,t)$, as suggested by Theorem 4. We denote the Levi-Civita connection of the metric $g_{0}$ by $D$. We then define
    $\bar{D}_{X}Y:=D_{X}Y+\hat{S}(X,Y),$
    for any $X,Y\in\mathfrak{X}(M)$, where $\hat{S}=g^{-1}S$. A short computation yields
    $\bar{D}g_{0}=-2(g_{0}\hat{S}),$
    since $Z(g_{0}(X,Y))=g_{0}(D_{Z}X,Y)+g_{0}(X,D_{Z}Y)$ for $X,Y,Z\in\mathfrak{X}(M)$. Thereby, the abundant structure $(M,g,S,t)$ gives rise to the semi-Weyl structure with data $(M,g_{0},\bar{D},0)$. More generally, we obtain the following theorem.
    Theorem 5.
    An abundant structure $(M,g,S,t)$ has an underlying semi-Weyl structure. In particular, denoting by $\nabla^{0}$ the Levi-Civita connection of $\exp(-\frac{2}{3}t)g$, let
    $\bar{D}:=\nabla^{0}+\hat{S}$
    where $\hat{S}=g^{-1}S$. For $h=\Omega^{2}g$ another metric in the same conformal class as $g$, it then follows that $(M,h,\bar{D},\theta:=-\frac{2}{3}dt-2d\ln|\Omega|)$ is a semi-Weyl structure, and
    $\bar{D}_{Z}h(X,Y)+\theta(Z)h(X,Y)\in\Gamma(\mathrm{Sym}^{3}(T^{*}M)).$
    Proof.
    Without loss of generality, let $g=g_{0}$, such that $h=\Omega^{2}g_{0}$ and $t=0$. A straightforward computation shows
    $\bar{D}_{Z}(\Omega^{2}g_{0})-2\left(d\ln|\Omega|\right)(Z)\,\Omega^{2}\,g_{0}(X,Y)\\ =2\Omega\,d\Omega(Z)g_{0}(X,Y)+\Omega^{2}\bar{D}_{Z}g_{0}(X,Y)-2\left(d\ln|\Omega|\right)(Z)\,\Omega^{2}\,g_{0}(X,Y)\\ =-2\Omega^{2}(g_{0}\hat{S})(X,Y,Z).$
    The general case is proven analogously. ∎
    We have hence found that the data of $(M,g,S,t)$ can alternatively be encoded in the semi-Weyl structure $(M,g,\bar{D},-\frac{2}{3}dt)$ where we identify abundant structures $(g,S,t)\sim(g,S,t^{\prime})$ when $t-t^{\prime}$ is a constant.
    We also mention that this endows $(M,g_{0},g_{0}g^{-1}S)$ with the structure of a statistical manifold, cf. [Matsuzoe2010].